# Cross-modal RAG: Sub-dimensional Retrieval-Augmented Text-to-Image Generation

### v2

@@ Appendix E Limitations
While our multi-objective joint retrieval combining sub-dimensional sparse and dense retrievers is efficient and achieves good granularity, when it comes to generation, the granularity of control is still bounded by the capabilities of the underlying MLLM. If the underlying MLLM lacks the granularity of control in generation, it becomes difficult to effectively leverage the subquery-level information obtained from our retrieval method.
Appendix F More Visualizations of Our Method Compared with Other Baselines
More visualizations of our proposed method Cross-modal RAG compared with other baselines can be found in Fig. 6 to  8. Across all three datasets, our method achieves superior image generation capability. For the CUB in Fig. 6, our Cross-modal RAG can generate realistic images that align with all subqueries, which are often ignored or distorted in GILL and UniRAG. Besides, SDXL, LaVIT, and RDM tend to generate the sketch-like images rather than photo-realistic birds. On the ImageNet-LT in Fig. 7, retrieving relevant images plays a crucial role in generating accurate long-tailed objects. Our method successfully generates all three long-tailed objects with high visual fidelity. In contrast, none of the baselines are able to generate all three correctly - UniRAG and GILL even fail to produce a single accurate image. For WikiArt in Fig. 8, creative image generation poses a unique challenge, as it is inherently difficult to reproduce the exact ground-truth image. However, our method explicitly retrieve the images with satisfied subqueries and can capture the particular artistic style specified in the query. As a result, all three generated images of Cross-modal RAG closely resemble the style of the target artist in the query. In contrast, other RAG baselines can not guarantee if the retrieved images grounded in the intended artist’s style. RDM even suffers from low visual fidelity when generating human faces.
Figure 6: Visualizations on CUB compared with other baselines.
Figure 7: Visualizations on ImageNet-LT compared with other baselines.
Figure 8: Visualizations on WikiArt compared with other baselines.

### v3

@@ Appendix E Limitations
While our multi-objective joint retrieval combining sub-dimensional sparse and dense retrievers is efficient and achieves good granularity, if the image retrieval database $\mathcal{D}$ does not contain any images relevant to the query, the retrieval cannot provide benefits to the generation. However, this drawback is not an issue unique to our RAG model; rather, the success of any RAG approach depends on the presence of relevant data (i.e., images in our case) in the database. As long as there exist partially related images in the database—even if they do not perfectly match all aspects of the query—our method can identify the most overlapping images with the query to effectively support generation.
Appendix F LLM Usage Disclosure
We use large language models to correct the grammar and improve the clarity of writing in this paper.
Appendix G Hyper-parameter analysis
For the trade-off parameter $\beta$ and the weight vector $\alpha$, in our experiments, we set $\beta=0.015$ based on Theorem 3.2, where for any $0<\beta<\beta_{\max}=\frac{\delta_{\min}}{C_{\max}}$, our algorithm guarantees to return all Pareto-optimal solutions. The maximum number of subqueries is $8$ in our experiments. Assuming equal weight for each subquery, we have $\beta_{\max}=\frac{1/8}{8}\approx 0.0156,$ so we set $\beta=0.015$ accordingly.
The design of $\alpha$ is intended to ensure the selection of a Pareto-optimal set of images would collectively cover all subqueries in the input user query. If an image satisfies a subquery that is not covered by other images, then there exists at least one choice of $\{\alpha_{i}\}$ for which it can maximize $\sum_{i=1}^{n}\alpha_{i}s_{i}(I_{j})$, allowing that image to be selected. Therefore, in our experiments, we vary $\alpha$ across all possible extreme values to ensure that every possible combination of subqueries is taken into account.
Appendix H More Visualizations of Our Method Compared with Other Baselines
More visualizations of our proposed method Cross-modal RAG compared with other baselines can be found in Fig. 6 to  8. Across all three datasets, our method achieves superior image generation capability. For the CUB in Fig. 6, our Cross-modal RAG can generate realistic images that align with all subqueries, which are often ignored or distorted in GILL and UniRAG. Besides, SDXL, LaVIT, and RDM tend to generate the sketch-like images rather than photo-realistic birds. On the ImageNet-LT in Fig. 7, retrieving relevant images plays a crucial role in generating accurate long-tailed objects. Our method successfully generates all three long-tailed objects with high visual fidelity. In contrast, none of the baselines are able to generate all three correctly - UniRAG and GILL even fail to produce a single accurate image. For WikiArt in Fig. 8, creative image generation poses a unique challenge, as it is inherently difficult to reproduce the exact ground-truth image. However, our method explicitly retrieve the images with satisfied subqueries and can capture the particular artistic style specified in the query. As a result, all six generated images of Cross-modal RAG closely resemble the style of the target artist in the query. In contrast, other RAG baselines can not guarantee if the retrieved images grounded in the intended artist’s style. RDM even suffers from low visual fidelity when generating human faces.
Figure 6: Visualizations of generation on CUB compared with other baselines.
Figure 7: Visualizations of generation on ImageNet-LT compared with other baselines.
Figure 8: Visualizations of generation on WikiArt compared with other baselines.